# Artin–Schreier quandles of involutions in absolute Galois groups

### 12 Mar 2024 v1

@@ 5 Laurent series and other examples
In this section, we are going to address two questions. First, for any given field $F$, we will describe the involutory quandle $\mathrm{AS}(F(\!(t)\!))$ of the field $F(\!(t)\!)$ of Laurent series with coefficients in $F$ in terms of the Artin–Schreier quandle $\mathrm{AS}(F)$. Second, we comment on the class of profinite involutory quandles that are realisable, up to isomorphism, as Artin–Schreier quandles $\mathrm{AS}(F)$ for some field $F$. From what we have seen for local and global fields so far, it seems reasonable to conjecture that these are all free, but we will show by example that this is not the case (see also our Propositions 3.11 and 3.17 for other restrictions).
@@
$\mathrm{AS}(F(\!(t)\!))\cong\mathrm{AS}(F)\times\mathrm{FQ}^{\wedge}_{2}\{\sigma,\tau\}$
of involutory quandles, where $\mathrm{FQ}^{\wedge}_{2}\{\sigma,\tau\}$ is the free pro-finite involutory quandle on a set $\{\sigma,\tau\}$ with two elements.
Before we embark on the proof, let us recall that we can identify $\mathrm{FQ}^{\wedge}_{2}\{\sigma,\tau\}$ with the set of involutions in the free pro-finite $2$-group $\mathrm{F}^{\wedge}_{2}\{\sigma,\tau\}$ on those generators, which is the pro-finite infinite dihedral group $\widehat{\mathrm{D}}_{\infty}$. From the extension ^$\widehat{\mathbb{Z}}\to\widehat{\mathrm{D}}_{\infty}\to\mathbb{Z}/2$ of groups, we see that we can also identify $\mathrm{FQ}^{\wedge}_{2}\{\sigma,\tau\}$ with the set $\widehat{\mathbb{Z}}$, where the quandle structure is given by $a\rhd b=2a-b$.
@@
The statement is void if the field $F$ is not formally real: if the Laurent series field $F(\!(t)\!)$ can be ordered, so can the subfield $F$. In particular, we can assume that the field $F$ has characteristic 00. We start by choosing an algebraic closure $F(\!(t)\!)^{\mathrm{alg}}$ of the Laurent series field $F(\!(t)\!)$. Inside it, we find the field $F^{\mathrm{alg}}(\!(t)\!)$, where $F^{\mathrm{alg}}$ is the algebraic closure of the field $F$ inside $F(\!(t)\!)^{\mathrm{alg}}$. Then, we have a short exact sequence
$1\longrightarrow\operatorname{Gal}(F^{\mathrm{alg}}(\!(t)\!))\longrightarrow\operatorname{Gal}(F(\!(t)\!))\longrightarrow\operatorname{Gal}(F)\longrightarrow 1$
of absolute Galois groups, where the automorphism group $\operatorname{Aut}(F^{\mathrm{alg}}(\!(t)\!)|F(\!(t)\!))$ has been identified with $\operatorname{Gal}(F)$. A splitting of this short exact sequence is given by letting $\operatorname{Gal}(F)$ act on the coefficients of the elements in $F(\!(t)\!)$. We deduce that the absolute Galois group $\operatorname{Gal}(F(\!(t)\!))$ is a semidirect product  (compare [14]). We know that the field $F^{\mathrm{alg}}(\!(t)\!)$ is quasi-finite, i.e., we have a continuous isomorphism $\operatorname{Gal}(F^{\mathrm{alg}}(\!(t)\!))\cong\widehat{\mathbb{Z}}$ of groups. Thus, we have an identification $\operatorname{Gal}(F^{\mathrm{alg}}(\!(t)\!))\cong\widehat{\mathbb{Z}}\rtimes\operatorname{Gal}(F)$, and the multiplication takes the form
$(a,g)(b,h)=(a+g(b),gh).$
In particular, an element $(a,g)$ is an involution if and only if $g=s$ is an involution and $a+s(a)=0$. We claim that every involution $s\in\operatorname{Gal}(F)$ acts as $-1$ on the kernel $\widehat{\mathbb{Z}}$. For, if $s$ acts as the identity, then $a+s(a)=0$ implies $02a=$ 0, hence $a=0$ because $\widehat{\mathbb{Z}}$ is torsionfree. In other words, the element $s$ in $\operatorname{Gal}(F)$ would have a unique pre-image in $\operatorname{Gal}(F^{\mathrm{alg}}(\!(t)\!))$, and the same would hold for the conjugates of $s$. But for every ordering of $F$, there are two different orderings of $F^{\mathrm{alg}}(\!(t)\!))$, one where $t$ is positive and one where $t$ is negative (see [32, Prop. VIII.4.11], for instance). Hence, this is not possible, and $s$ must act as $-1$. But when any involution $s$ acts as $-1$, the equation $a+s(a)=0$ is satisfied by all elements in $\widehat{\mathbb{Z}}$, and we can identify the pre-image $\{(a,s)\,|\,a\in\widehat{\mathbb{Z}}\}$ of any involution $s$ canonically with $\widehat{\mathbb{Z}}$. Taken together, we get a bijection between $AS(F(\!(t)\!))$ and the product $\mathrm{AS}(F)\times\widehat{\mathbb{Z}}$. The quandle operation is induced from the conjugation in the group, and the computation
@@
We close by mentioning some other examples that show similar features to the above.
@@
As the Artin–Schreier quandle $\mathrm{AS}(\mathbb{R})$ is a singleton, Theorem 5.1 immediately gives that $\mathrm{AS}(\mathbb{R}(\!(t)\!))$ is free, isomorphic to $\mathrm{FQ}^{\wedge}_{2}\{\sigma,\tau\}$. By intersecting two real closed fields, Geyer [22] has constructed other formally real fields whose absolute Galois group is infinite dihedral, i.e., isomorphic to the free pro-2-group $\mathrm{F}^{\wedge}_{2}\{\sigma,\tau\}$ on two generators. Of course, the correspoding Artin–Schreier quandle $\mathrm{FQ}^{\wedge}_{2}\{\sigma,\tau\}$ is free on two generators, too. However, applying Theorem 5.1 twice, we see that
@@
and I see no reason why this pro-finite involutory quandle should be free.
Example 5.3.
Let $R$ be the real closure of $\mathbb{Q}$ inside $\mathbb{R}$. Consider subfields of $R$ such that do not contain a square root of $2$, and assume that $F$ is maximal among those. Then the absolute Galois group $\operatorname{Gal}(F)$ is the semi-direct product $\mathrm{F}^{\wedge}_{2}\{\rho\}\ltimes\mathrm{F}^{\wedge}_{2}\{\sigma,\tau\}$, where $\rho$ acts by conjugating the other two generators into each other. We refer to Engler–Viswanathan [16] for details. It is clear, though, that the relation $\rho\rhd\sigma=\tau$ is fulfilled in $\mathrm{AS}(F)$, which is, therefore, not free.
@@ Acknowledgements
I thank Jayanta Manoharmayum for discussions about absolute Galois groups.

### 08 Apr 2026 v2

@@ 5 Laurent series and other examples
In this section, we will address two questions. First, for any given field $F$, we will describe the involutory quandle $\mathrm{AS}(F(\!(t)\!))$ of the field F((t))F(\!(t)\!) of Laurent series with coefficients in $F$ in terms of the Artin–Schreier quandle $\mathrm{AS}(F)$. Second, we comment on the class of profinite involutory quandles that are realisable, up to isomorphism, as Artin–Schreier quandles $\mathrm{AS}(F)$ for some field $F$. From what we have seen for local and global fields so far, it seems reasonable to conjecture that these are all free, but we will show by example that this is not the case (see also our Propositions 3.11 and 3.17 for other restrictions).
@@
$\mathrm{AS}(F(\!(t)\!))\cong\mathrm{AS}(F)\times\mathrm{FQ}^{\wedge}_{2}\{\sigma,\tau\}\cong\mathrm{AS}(F)\times\widehat{\mathbb{Z}}$
of profinite involutory quandles, where $\mathrm{FQ}^{\wedge}_{2}\{\sigma,\tau\}$ is the free profinite involutory quandle on a set $\{\sigma,\tau\}$ with two elements. The quandle structure is given by $a\rhd b=2a-b$ for $a,b\in\widehat{\mathbb{Z}}$.
Before we embark on the proof, let us recall that we can identify $\mathrm{FQ}^{\wedge}_{2}\{\sigma,\tau\}$ with the set of involutions in the free profinite 22–group $\mathrm{F}^{\wedge}_{2}\{\sigma,\tau\}$. This group $\mathrm{F}^{\wedge}_{2}\{\sigma,\tau\}$ is the profinite infinite dihedral group $\widehat{\mathrm{D}}_{\infty}$, and we can identify it with the semi-direct product $\widehat{\mathbb{Z}}\rtimes\mathbb{Z}/2$, where the generator of $\mathbb{Z}/2$ acts as 1-1 on $\widehat{\mathbb{Z}}$. An isomorphism is given by letting a pair $(a,m)$ in the semi-direct product $\widehat{\mathbb{Z}}\rtimes\mathbb{Z}/2$ correspond to $(\sigma\tau)^{a}\sigma^{m}$ in $\mathrm{F}^{\wedge}_{2}\{\sigma,\tau\}$. Once the groups are identified this way, we see that we can thereby also identify the subset $\mathrm{FQ}^{\wedge}_{2}\{\sigma,\tau\}$ of involutions with the set $\widehat{\mathbb{Z}}$, where now $a\in\widehat{\mathbb{Z}}$ corresponds to the pair $(a,m)$ with $m$ odd in $\widehat{\mathbb{Z}}\rtimes\mathbb{Z}/2$. For $m$ and $n$ odd, we get
$(a,m)(b,n)(a,m)^{-1}$ $=(a+(-1)^{m}b,m+n)((-1)^{-m}(-a),-m)$
$=(a+(-1)^{m}b+(-1)^{m+n}(-1)^{-m}(-a),m+n-m)$
$=(2a-b,n),$
and this calculation shows that the quandle structure we get on the set $\widehat{\mathbb{Z}}$ by transport of structure is given by $a\rhd b=2a-b$.
@@
The statement is void if the field $F$ is not formally real: if the Laurent series field F((t))F(\!(t)\!) can be ordered, so can the subfield $F$. In particular, we can assume that the field $F$ has characteristic 0. For these fields, it is known that the absolute Galois group $\operatorname{Gal}(F(\!(t)\!))$ is a semidirect product (cf. [16]); in fact, there is an isomorphism $\operatorname{Gal}(F(\!(t)\!))\cong\widehat{\mathbb{Z}}\rtimes\operatorname{Gal}(F)$ (see [23, Prop. 4.1(e)]). The multiplication takes the form
$(a,g)(b,h)=(a+g(b),gh)$
for some $\operatorname{Gal}(F)$–action on $\widehat{\mathbb{Z}}$. In particular, an element $(a,g)$ is an involution if and only if $g=s$ is an involution and $a+s(a)=0$.
If an involution $s$ acted as the identity, then $a+s(a)=0$ implies $2a=0$, hence $a=0$ because $\widehat{\mathbb{Z}}$ is torsionfree. In other words, the element $s$ in $\mathrm{AS}(F)=\operatorname{Inv}(\operatorname{Gal}(F))$ would have a unique pre-image in $\mathrm{AS}(F(\!(t)\!))$, and the same would hold for the conjugates of $s$. But for every ordering of $F$, there are two different orderings of F((t))F(\!(t)\!), one where $t$ is positive and one where $t$ is negative (see [35, Prop. VIII.4.11], for instance). Hence, this is not possible, and $s$ must act non-trivially. In fact, we can be more precise:
Lemma 5.2.
Every involution $s\in\operatorname{Gal}(F)$ acts as 1-1 on the kernel $\widehat{\mathbb{Z}}$.
If we had an action on the group $\mathbb{Z}$, it would be clear that an involution $s$ can only act as $\pm 1$, but $\widehat{\mathbb{Z}}$ has more units than these, and we have to argue to exclude any other options. We will do this in detail once we have finished the main argument.
When the involution $s$ acts as 1-1, the equation $a+s(a)=0$ is satisfied by all elements in $\widehat{\mathbb{Z}}$, and we can identify the pre-image $\{(a,s)\mid a\in\widehat{\mathbb{Z}}\}$ of any involution $s$ canonically with $\widehat{\mathbb{Z}}$. Taken together, we thereby get a bijection between $\mathrm{AS}(F(\!(t)\!))$ and the product $\mathrm{AS}(F)\times\widehat{\mathbb{Z}}$. The quandle operation is induced from the conjugation in the group, and the computation
@@
Proof of Lemma 5.2.
Let $\overline{F}$ be our chosen algebraic closure of $F$. Puiseux’s theorem states that the algebraic closure of $\overline{F}(\!(t)\!)$ is the union $\cup_{n}\overline{F}(\!(t^{1/n})\!)$, which we will denote by $\overline{F}(\!(t^{1/\infty})\!)$. For the field F((t))F(\!(t)\!), we need to be a bit more careful (see [23, Prop. 4.1] again). An algebraic closure of F((t))F(\!(t)\!) is the compositum $\overline{F}\cdot F(\!(t^{1/\infty})\!)$. It contains the subfields $F(\!(t^{1/\infty})\!)$ and $\overline{F}\cdot F(\!(t)\!)$, with absolute Galois groups isomorphic to $\operatorname{Gal}(F)$ and $\widehat{\mathbb{Z}}$, respectively. The latter is normal, and we find that $\operatorname{Gal}(F(\!(t)\!))\cong\widehat{\mathbb{Z}}\rtimes\operatorname{Gal}(F)$ is a semi-direct product.
To better understand the $\operatorname{Gal}(F)$–action on $\widehat{\mathbb{Z}}$, we proceed as follows. A generator $F_{n}$ of $\operatorname{Gal}(\overline{F}\cdot F(\!(t^{1/n})\!)\mid\overline{F}\cdot F(\!(t)\!))\cong\mathbb{Z}/n$ acts by $F_{n}(t^{1/n})=w_{n}t^{1/n}$ for a primitive nn–th root of unity $w_{n}\in\mu_{n}(\overline{F})$. The elements $F_{n}$ and $w_{n}$ can be chosen with $(w_{mn})^{m}=w_{n}$ for all $m$ so that the generators $F_{n}$ fit together to define a topological generator of $\operatorname{Gal}(\overline{F}\cdot F(\!(t)\!))\cong\widehat{\mathbb{Z}}$. The absolute Galois group $\operatorname{Gal}(F)$ acts on the coefficients of the Laurent series. On the normal subgroup $\operatorname{Gal}(\overline{F}\cdot F(\!(t)\!))$ of $\operatorname{Gal}(F(\!(t)\!))$, it acts by conjugation, but the computation
$sF_{n}s^{-1}(t^{1/n})=sF_{n}(t^{1/n})=s(w_{n}t^{1/n})=s(w_{n})t^{1/n}$
shows that this conjugation action is also just an action on the coefficients. In summary, we see that the $\operatorname{Gal}(F)$–action on $\widehat{\mathbb{Z}}$ is determined by the $\operatorname{Gal}(F)$–action on the group $\mu_{\infty}(\overline{F})$ of roots of unity in the algebraic closure $\overline{F}$.
It remains for us to check that all involutions in the absolute Galois group $\operatorname{Gal}(F)$ act on the group $\mu_{\infty}(\overline{F})$ of roots of unity in the algebraic closure $\overline{F}$ via the inversion $w\mapsto w^{-1}$. We argue one prime at a time. If the prime $p$ is odd, then all primitive pp–th power roots of unity are totally imaginary, and they cannot be fixed by an involution $s$. As there is only one involution in the cyclic group $(\mathbb{Z}/p^{e})^{\times}$, this forces $s(w)=w^{-1}$ for all $w\in\mu_{p^{e}}(\overline{F})$. At the even prime $2$, when $w\in\mu_{2^{e}}(\overline{F})$, we must have $s(w)=w^{-1}$ for $e=1$ trivially and for $e=2$ by a similar reasoning as above: the group $(\mathbb{Z}/4)^{\times}$ is cyclic. However, for $e\geqslant 3$, there are two other options for us to consider, namely s(w)=ws(w)=-w and $s(w)=-w^{-1}$. The first would imply that we have $s(w^{2})=w^{2}$, which we can rule out by induction on $e$, as $w^{2}\in\mu_{2^{e-1}}(\overline{F})$. The second would imply that $w-w^{-1}$ were fixed by $s$, but this element is totally imaginary. That leaves us with $s(w)=w^{-1}$ for all $w\in\mu_{2^{e}}(\overline{F})$ as well. ∎
@@
As the Artin–Schreier quandle $\mathrm{AS}(\mathbb{R})$ is a singleton, Theorem 5.1 immediately gives that the Artin–Schreier quandle $\mathrm{AS}(\mathbb{R}(\!(t)\!))$ is free, isomorphic to $\mathrm{FQ}^{\wedge}_{2}\{\sigma,\tau\}$. By intersecting two real closed fields, Geyer [22] has constructed other formally real fields whose absolute Galois group is infinite dihedral, i.e., isomorphic to the free pro-2-group $\mathrm{F}^{\wedge}_{2}\{\sigma,\tau\}$ on two generators. Of course, the correspoding Artin–Schreier quandle $\mathrm{FQ}^{\wedge}_{2}\{\sigma,\tau\}$ is free on two generators as well.
We conclude with an example showing that Artin–Schreier quandles need not be free.
Example 5.4.
Applying Theorem 5.1 twice, we see that
@@
and this profinite involutory quandle is not free. In fact, it is not even isomorphic to a quandle of the form $\mathrm{FQ}^{\wedge}_{2}(S)$ after forgetting the topologies, i.e., algebraically. To see this, we can consider the group of transvections (or displacements) of these quandles, introduced by Joyce [30, Sec. 5]: for a given quandle $Q$, the subgroup $\operatorname{Tr}(Q)\leqslant\operatorname{Aut}(Q)$ is the subgroup generated by the differences ${\lambda_{a}}^{\!\!\!-1}\lambda_{b}$ for all $a,b\in Q$. If $S$ has at most one element, this group is trivial. As $\mathrm{FQ}^{\wedge}_{2}\{\sigma,\tau\}\cong\widehat{\mathbb{Z}}$ with $a\rhd x=2a-x$, we can identify the group $\operatorname{Tr}(\mathrm{FQ}^{\wedge}_{2}\{\sigma,\tau\})$ with the group of maps $\widehat{\mathbb{Z}}\to\widehat{\mathbb{Z}}$ of the form $x\mapsto x+2c$, and this group is isomorphic to $\widehat{\mathbb{Z}}$ again. As $\operatorname{Tr}(Q\times R)\cong\operatorname{Tr}(Q)\times\operatorname{Tr}(R)$ by [31, Lem. 5.1], we have $\operatorname{Tr}(\mathrm{AS}(\mathbb{R}(\!(t_{1})\!)(\!(t_{2})\!)))\cong\widehat{\mathbb{Z}}^{2}$ for the quandle under consideration. It remains for us to argue that the groups $\operatorname{Tr}(\mathrm{FQ}^{\wedge}_{2}(S))$ are never of this form. From what we have computed above, this is already clear for $|S|<3$. But if $S$ contains three different elements sstt, and $u$, the group $\operatorname{Tr}(\mathrm{FQ}^{\wedge}_{2}(S))$ is not even abelian, as the elements ${\lambda_{s}}^{\!\!\!-1}\lambda_{u}$ and ${\lambda_{t}}^{\!\!\!-1}\lambda_{u}$ do not commute (check on $u$).
@@ Acknowledgements
I thank Jayanta Manoharmayum for helpful discussions on absolute Galois groups, and the anonymous referee for exceptionally detailed and constructive comments that led to improvements in several aspects of this paper.